# Recommenadation aided Caching using Combinatorial Multi-armed Bandits

### v1

@@ II-A Content Caching Problem
In Problem (4), user preference distributions in the absence of recommendation, $p_{u}^{\text{pref}}(i)$ for all $u,i$, are unknown. One way of solving this problem is to first estimate $p_{u}^{\text{pref}}(i)$ for all $u,i$, and then solve the optimization problem. But recommendations affect the number of requests, so we cannot directly estimate $p_{u}^{\text{pref}}(i)$. We can only estimate the request distributions $p_{u}^{\text{req}}(i)$. Also, we can only observe the requests of the cached contents and hence it is modeled as CMAB(i.e., in CMAB , only the rewards of the pulled arms are observed). This is different from full observation setting [1] where we can observe the requests of all the contents irrespective of whether the content is cached or not. So in the next section, we introduce a UCB-based algorithm that estimates $p_{u}^{\text{req}}(i)$ for all $u,i$, and solves Problem (4).
@@ III Algorithm Design
If we use Algorithm 1 without recommendation, i.e., setting $\bar{w}_{u}^{\text{rec}}=0$ for all $u$, then we get the standard CMAB. An upper bound on its regret can be derived following similar approach as in [10, Theorem 2]. It turns out to be as follows.
@@
The proof of Theorem 1 is motivated by [10, Theorem 2]. However, we use different confidence intervals, which warrant different arguments at several steps.
@@
${\leqslant}\frac{16\left(1-\bar{w}^{\text{rec}}\right)^{1/\eta}\log T}{\Delta_{\min,k}}+\frac{4\alpha\left(\bar{w}^{\text{rec}}\right)}{\Delta_{\min,k}},$ (23)
@@ IV Numerical Results
We compare the performance of Algorithm (1) with CMAB-UCB with $\epsilon$-Greedy and Greedy algorithm. In the Greedy algorithm, we calculate $p_{i}(t)$ for all $i\in\mathcal{N}$. Then, pick the top $C$ contents with the highest $p_{i}(t)$. In the $\epsilon$-Greedy algorithm, at each time $t$, a random number is generated every time. If it exceeds the epsilon, the top $C$ contents with the highest $p_{i}(t)$ are selected. Else $C$ contents selected uniformly from $\mathcal{N}=\{1,2,3,...,N\}$. In this experiment, we used $w_{u}^{rec}=0.95$ for all $u\in\mathcal{U}$, $\alpha=5,\epsilon=0.4$. We can see from the figure Fig.(1) that our algorithm performs better compared to better to other algorithms.
@@
In this experiment will compare the performance of our algorithm for various $w_{u}^{rec}$ values i.e. $w_{u}^{rec}=0.99,0.95,0.9$ for all $u\in\mathcal{U}$. When $w_{u}^{rec}\approx 1$, users always tend to accept the recommendation. Since we recommending whatever is cached, we expect the regret to be low. Hence, we want regret to decrease when $w_{u}^{rec}$ increases and becomes close to 1. This can be seen from the Fig.(2)
@@ V Conclusion and Future Work
In this work we analysed the content caching problem with recommendation, modeling it as a CMAB. We proposed an UCB-based algortihm and also provided an upper bound on the regret of the algorithm. We then compared the performance of this algorithm to that of a few existing algorithms. We also considered the case where the position of the recommended content affects its request probability. We found that caching with recommendation improves the cache hit performance in all the cases.
In our future work, we would like to consider a different recommendation model where each user examines the contents in the recommendation list one by one until he finds attractive content and does not examine the rest. We would like to incorporate recommendation quality (13) into our problem and would also like to consider the case where $w_{u}^{\mathrm{rec}}$’s are unknown.

### v2

# Recommendation aided Caching using Combinatorial Multi-armed Bandits

@@ II-A Content Caching Problem
In Problem (4), user preference distributions in the absence of recommendation, $p_{u}^{\text{pref}}(i)$ for all $u,i$, are unknown. One way of solving this problem is to first estimate $p_{u}^{\text{pref}}(i)$ for all $u,i$, and then solve the optimization problem. But recommendations affect the number of requests, so we cannot directly estimate $p_{u}^{\text{pref}}(i)$. We can only estimate the request distributions $p_{u}^{\text{req}}(i)$. Also, we can only observe the requests of the cached contents and hence it is modeled as CMAB(i.e., in CMAB , only the rewards of the pulled arms are observed). This is different from full observation setting where we can observe the requests of all the contents irrespective of whether the content is cached or not. So in the next section, we introduce a UCB-based algorithm that estimates $p_{u}^{\text{req}}(i)$ for all $u,i$, and solves Problem (4).
@@ III Algorithm Design
If we use Algorithm 1 without recommendation, i.e., setting $\bar{w}_{u}^{\text{rec}}=0$ for all $u$, then we get the standard CMAB. An upper bound on its regret can be derived following similar approach as in [9, Theorem 2]. It turns out to be as follows.
@@
The proof of Theorem 1 is motivated by [9, Theorem 2]. However, we use different confidence intervals, which warrant different arguments at several steps.
@@
${\leqslant}\frac{16\left(1-\bar{w}^{\text{rec}}\right)^{1/\eta}\log T}{\Delta_{\min,k}}+\frac{4\alpha\left(\bar{w}^{\text{rec}}\right)}{\Delta_{\min,k}}$ (23)
@@ IV Numerical Results
We compare the performance of Algorithm (1) with CMAB-UCB ,$\epsilon$-Greedy and Greedy algorithm. In the Greedy algorithm, we calculate $p_{i}(t)$ for all $i\in\mathcal{N}$. Then, pick the top $C$ contents with the highest $p_{i}(t)$. In the $\epsilon$-Greedy algorithm, a random number is generated every time. If it exceeds the epsilon, the top $C$ contents with the highest $p_{i}(t)$ are selected. Else $C$ contents selected uniformly from content set $\mathcal{N}$. In this experiment, we used $w_{u}^{rec}=0.95$ for all $u\in\mathcal{U}$, $\alpha=5,\epsilon=0.4$. We can see from the figure Fig.(1) that our algorithm performs better compared to better to other algorithms.
@@
In this experiment will compare the performance of our algorithm for various $w_{u}^{rec}$ values i.e. $w_{u}^{rec}=0.99,0.95,0.9$ for all $u\in\mathcal{U}$. When $w_{u}^{rec}\approx 1$, users always tend to accept the recommendation. Since we recommending cache contents, we expect the regret to be low. Hence, we want regret to decrease when $w_{u}^{rec}$ increases and becomes close to 1. This can be seen from the Fig.(2)
@@ V Conclusion and Future Work
In this work, we analysed the content caching problem with the recommendations, modelling it as a CMAB. We proposed a UCB-based algorithm and also provided an upper bound on the regret of the algorithm. We then compared the performance of this algorithm to that of a few existing algorithms. We also considered the case where the position of the recommended content affects its request probability. We found that caching with recommendation improves the cache hit performance in all the cases.
In our future work, we would like to consider a different recommendation model where each user examines the contents in the recommendation list one by one until he finds attractive content, and once he finds any attractive content, he does not examine the rest. We would like to incorporate recommendation quality (13) into our problem and would also like to consider the case where $w_{u}^{\mathrm{rec}}$’s are unknown.